OpenStudy (anonymous):

x^2=-2x+33 factor show steps

OpenStudy (anonymous):

x^2+2x-33 now use the ac method and i will check your work

OpenStudy (anonymous):

x2=−2x+33 x2+2x=−2x+33+2x x2+2x=33 x2+2x+1=33+1 x2+2x+1=34 ?

OpenStudy (usukidoll):

it's correct so far... op is using completing the square

OpenStudy (anonymous):

what do u mean

OpenStudy (usukidoll):

completing a square is a method used to solve a quadratic equation.

OpenStudy (anonymous):

im trying to factor

OpenStudy (usukidoll):

this is not factoring.. this is solving an equation with completing the square. you're not done yet. What you need to do is to convert the left hand side into squared form

OpenStudy (anonymous):

how do i factor, i need to factor it

OpenStudy (usukidoll):

x^2+2x-33=0 is impossible to factor.. 33 is too big! all combinations of 33 is too big. You either need to use the quadratic formula or complete the square.

OpenStudy (usukidoll):

3 x 11 = 33 3+11 = 14 or 3-11 = -8 33 x 1 = 33 33+1 = 34 or 33-1 = 32 having a 2 to show up is impossible!

OpenStudy (anonymous):

WAIT THIS IS FROM MY TEST

OpenStudy (anonymous):

._.

OpenStudy (usukidoll):

huh? ok what is going on?

OpenStudy (anonymous):

this is a math problem from my test you write down ur answer with all the steps and take a picture of it then send it to the teacher

OpenStudy (usukidoll):

oooohhhhhh... you do realize that posting a question from a quiz/exam/test is against Openstudy's rules right?

OpenStudy (anonymous):

its homework tho

OpenStudy (usukidoll):

so where did this come from? someone's test or something else?

OpenStudy (anonymous):

lol a connexus test bai